# Hierarchical Auto-Organizing System for Open-Ended Multi-Agent Navigation

### v1

@@ Abstract
Navigating complex environments in Minecraft poses significant challenges for multi-agent systems due to the game’s dynamic and unpredictable open-world setting. Agents need to interact with the environment and coordinate their actions with other agents to achieve common objectives. However, traditional approaches often struggle to efficiently manage inter-agent communication and task distribution, which are crucial for effective multi-agent navigation. Furthermore, processing and integrating multi-modal information (such as visual, textual, and auditory data) is essential for agents to fully comprehend their goals and navigate the environment successfully. To address this issue, we design the HAS framework to auto-organize groups of LLM-based agents to complete Navigation tasks. In our approach, we devise a hierarchical auto-organizing navigation system, which is characterized by 1) a hierarchical system for multi-agent organization, ensuring centralized planning and decentralized execution; 2) an auto-organizing and intra-communication mechanism, enabling dynamic group adjustment under subtasks; 3) a multi-modal information platform, facilitating multi-modal perception to perform the three navigation tasks with one system. To assess organizational behavior, we design a series of navigation tasks in the Minecraft environment, which includes searching and exploring. We aim to develop embodied organizations that push the boundaries of embodied AI, moving it towards a more human-like organizational structure.
@@ 1 Introduction
Drawing on large language models (LLMs) with human knowledge, communicative competence, and decision-making capabilities, embodied agents exhibit human-like intelligence in playing games (Park et al., 2023; Wang et al., 2023a; Zhao et al., 2023), programming (Qian et al., 2023; Hong et al., 2023), and robotic tasks (Zhang et al., 2023; Mandi et al., 2023). The development of individual intelligence has led to creating a new, collaborative framework where multiple agents Chen et al. (2023a; b) work together.
This shift towards a multi-agent system Hong et al. (2023); Chen et al. (2023a) uses advanced language understanding and decision-making abilities to enhance intelligence through interactions and data sharing. Agents in this system specialize in various tasks, sharing insights to enhance overall efficiency. This collaborative approach not only optimizes execution but also fosters a learning environment where agents continually improve through shared intelligence. Their interaction leads to refined skills, equipping them to handle complex, large-scale tasks through parallel processing and coordinated collaboration—a notable advancement in autonomous systems.
Multi-modal navigation (Wijmans et al., 2019; Yu et al., 2021; Du et al., 2020; Kwon et al., 2023; Chen et al., 2021; Moudgil et al., 2021) stands at the forefront of contemporary AI research, representing a rapidly evolving field that aims to integrate various sensory inputs into a cohesive navigational strategy. This emerging domain extends the capabilities of multi-agent systems by requiring them to interpret and act upon a confluence of sensory data. Unlike traditional approaches focusing solely on rendered images or static virtual environments, multi-modal navigation within open-ended environments like Minecraft presents a more complex and nuanced challenge. Agents in Minecraft Wang et al. (2023a); Zhao et al. (2023) navigate a world full of freedom and variability, making it an ideal testbed for embodied agent systems. Minecraft’s unpredictable and open-ended environment is an ideal testing ground for advanced AI systems. These systems emulate human adaptability and intelligence, processing and sharing multi-modal data to strategically navigate and complete tasks. This approach goes beyond simple improvements, marking a significant leap in developing autonomous systems that handle complex tasks with remarkable autonomy and skill.
Within this context, integrating multiple Multi-modal Language Models (MLMs) into embodied agents opens up new avenues for enhancing navigational efficiency. These agents, equipped with LLMs, are not just reactive but proactive, possessing the ability to engage with their environment intelligently and contextually. The zero-shot performance capabilities of LLMs, where they can interpret and act upon data without prior explicit programming, make them particularly suited for the dynamic challenges presented by open-ended environments.
@@ 2.3 LLM-based Multi-Agent Frameworks
Large Language Models (LLMs) are skilled at completing new tasks when given prompt-based instructions. Autonomous agents based on Large Language Model-based (LLM-based) models have gained significant interest in industry and academia (Wang et al., 2023b). Several works (Wang et al., 2023d; Du et al., 2023; Zhuge et al., 2023; Hao et al., 2023; Akata et al., 2023) have augmented the problem-solving abilities of LLMs by incorporating discussions among multiple agents. Stable-Alignment (Liu et al., 2023b) generates instruction datasets by reaching a consensus on value judgments through interactions among LLM agents in a sandbox. Some works in the field of artificial intelligence focus on studying sociological phenomena. For instance, Generative Agents (Park et al., 2023) creates a virtual ”town” comprising 25 agents to investigate language interaction, social understanding, and collective memory. The Natural Language-Based Society of Mind (Zhuge et al., 2023) involves agents with different functions interacting to solve complex tasks through multiple rounds of Mindstorms. In addition, others (Cai et al., 2023) propose a model for cost reduction by combining large models as tool makers and small models as tool users.
@@ 3.1 Centralized Planning with Decentralized Execution
Evoked from the Centralized Training with Decentralized Execution (CTDE) framework (Hu et al., 2023) for cooperative Multi-Agent Reinforcement Learning (MARL), we propose Centralized Planning with Decentralized Execution (CPDE) for the cooperative LLM-based Multi-Agent system. This architecture leverages global state information to inform the planning process, while the execution of tasks is carried out by local agents independently. CPDE is tailored to maximize the synergy of global oversight and local autonomy, ensuring efficient navigation and task completion in complex environments.
@@ 3.3 Multi-modal Memory
Retrival-augmented storage (RAS) enables long-term planning capability by Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Mao et al., 2020). RAG improves the quality of responses generated by language models with external sources of knowledge to complement the model’s internal representation. Instead of external knowledge libraries, we use the collected multi-modal memory as the knowledge library and retrieve interactive experiences as demonstration prompts to improve the planning results. The formulation is as follows:

### v2

@@ Abstract
Due to the dynamic and unpredictable open-world setting, navigating complex environments in Minecraft poses significant challenges for multi-agent systems. Agents must interact with the environment and coordinate their actions with other agents to achieve common objectives. However, traditional approaches often struggle to efficiently manage inter-agent communication and task distribution, crucial for effective multi-agent navigation. Furthermore, processing and integrating multi-modal information (such as visual, textual, and auditory data) is essential for agents to comprehend their goals and navigate the environment successfully and fully. To address this issue, we design the HAS framework to auto-organize groups of LLM-based agents to complete navigation tasks. In our approach, we devise a hierarchical auto-organizing navigation system, which is characterized by 1) a hierarchical system for multi-agent organization, ensuring centralized planning and decentralized execution; 2) an auto-organizing and intra-communication mechanism, enabling dynamic group adjustment under subtasks; 3) a multi-modal information platform, facilitating multi-modal perception to perform the three navigation tasks with one system. To assess organizational behavior, we design a series of navigation tasks in the Minecraft environment, which includes searching and exploring. We aim to develop embodied organizations that push the boundaries of embodied AI, moving it towards a more human-like organizational structure.
This work is supported by the National Natural Science Foundation of China (No. 62106219, No. 62206046), the Zhejiang Provincial Natural Science Foundation of China (No. LZ24F030005), and the Shanghai Sailing Program (No. 21YF1401300).
@@ 1 Introduction
Drawing on large language models (LLMs) with human knowledge, communicative competence Mao et al. (2020a; a), and decision-making capabilities, embodied agents exhibit human-like intelligence in playing games (Park et al., 2023; Wang et al., 2023a; Zhao et al., 2023), programming (Qian et al., 2023; Hong et al., 2023), and robotic tasks (Zhang et al., 2023b; Mandi et al., 2023). The development of individual intelligence has led to creating a new, collaborative framework where multiple agents Chen et al. (2023a; b) work together.
This shift towards a multi-agent system Hong et al. (2023); Chen et al. (2023a) uses advanced language understanding and decision-making abilities to enhance intelligence through interactions and data sharing. Agents in this system specialize in various tasks, sharing insights to enhance overall efficiency. This collaborative approach not only optimizes execution but also fosters a learning environment where agents continually improve through shared intelligence. Their interaction leads to refined skills, equipping them to handle complex, large-scale tasks through parallel processing and coordinated collaboration, a notable advancement in autonomous systems.
Multi-modal navigation (Wijmans et al., 2019; Yu et al., 2021; Du et al., 2020; Kwon et al., 2023; Chen et al., 2021; Moudgil et al., 2021) stands at the forefront of contemporary AI research, representing a rapidly evolving field that aims to integrate various sensory inputs into a cohesive navigational strategy. This emerging domain extends the capabilities of multi-agent systems by requiring them to interpret and act upon a confluence of sensory data. Unlike traditional methods focusing on rendered images or static virtual environments, works on dynamic environments Deng et al. (2023), such as multi-modal navigation in open-ended environments like Minecraft, pose a more complex challenge. Agents in Minecraft Wang et al. (2023a); Zhao et al. (2023) navigate a world full of freedom and variability, making it an ideal testbed for embodied agent systems. Minecraft’s unpredictable and open-ended environment is an ideal testing ground for advanced AI systems. These systems emulate human adaptability and intelligence, processing and sharing multi-modal data to strategically navigate and complete tasks. This approach goes beyond simple improvements, marking a significant leap in developing autonomous systems that handle complex tasks with remarkable autonomy and skill. Integrating Multi-modal Language Models (MLMs) into embodied agents enhances navigational efficiency. Equipped with LLMs, these agents engage intelligently with their environment. The zero-shot performance capabilities of LLMs enable them to interpret and act upon data without explicit programming. This makes them ideal for open-ended environments.
@@ 2.3 LLM-based Multi-Agent Frameworks
Large Language Models (LLMs) are skilled at completing new tasks when given prompt-based instructions. Autonomous agents based on Large Language Model-based (LLM-based) models have gained significant interest in industry and academia (Wang et al., 2023b). Several works (Wang et al., 2023d; Du et al., 2023; Zhuge et al., 2023; Hao et al., 2023; Akata et al., 2023; Zhang et al., 2023a) have augmented the problem-solving abilities of LLMs by incorporating discussions among multiple agents. Stable-Alignment (Liu et al., 2023b) generates instruction datasets by reaching a consensus on value judgments through interactions among LLM agents in a sandbox. Some works in the field of artificial intelligence focus on studying sociological phenomena. For instance, Generative Agents (Park et al., 2023) creates a virtual “town” comprising 25 agents to investigate language interaction, social understanding, and collective memory. The Natural Language-Based Society of Mind (Zhuge et al., 2023) involves agents with different functions interacting to solve complex tasks through multiple rounds of Mindstorms. In addition, others (Cai et al., 2023) propose a model for cost reduction by combining large models as tool makers and small models as tool users.
@@ 3.1 Centralized Planning with Decentralized Execution
Evoked from the Centralized Training with Decentralized Execution (CTDE) framework (Hu et al., 2023; Sunehag et al., 2017; Lowe et al., 2017; Mao et al., 2018) for cooperative Multi-Agent Reinforcement Learning (MARL), we propose Centralized Planning with Decentralized Execution (CPDE) for the cooperative LLM-based Multi-Agent system. This architecture leverages global state information to inform the planning process, while the execution of tasks is carried out by local agents independently. CPDE is tailored to maximize the synergy of global oversight and local autonomy, ensuring efficient navigation and task completion in complex environments.
@@ 3.3 Multi-modal Memory
Retrival-augmented storage (RAS) enables long-term planning capability by Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Mao et al., 2020b). RAG improves the quality of responses generated by language models with external sources of knowledge to complement the model’s internal representation. Instead of external knowledge libraries, we use the collected multi-modal memory as the knowledge library and retrieve interactive experiences as demonstration prompts to improve the planning results. The formulation is as follows: